OpenStudy (sugarmochi):

simplify (1-3n)^2 - (1+2n)(1-3n)

OpenStudy (jhannybean):

\[(a-b)(a+b) = a^2 -b^2\]\[(a-b)^2 = (a-b)^2-2ab\]

OpenStudy (jhannybean):

Use these two properties in helping you simplify your expression.

OpenStudy (butterflydreamer):

you could also factorise this question :). So look at both (1 - 3n)^2 and (1+2n) (1 - 3n) And factorise out what is common.

OpenStudy (sugarmochi):

the (1-3n)? ?

OpenStudy (butterflydreamer):

yaaa :D So factorise out " (1 - 3n) from the equation " (1 - 3n) ^2 - (1 + 2n)(1 - 3n)"

OpenStudy (sugarmochi):

so it'll become (1-3n) - (1+2n)?

OpenStudy (butterflydreamer):

hmm you're close but since we factorise out (1 - 3n) this means your equation actually looks like this: (1 - 3n)^2 - (1 + 2n)(1 - 3n) Since we factorised out (1 - 3n). We actually get: (1 - 3n) [ (1- 3n) - (1 + 2n)] Now you want to simplify what [ (1-3n) - (1+2n)]

OpenStudy (butterflydreamer):